# V2X-Radar: A Multi-modal Dataset with 4D Radar for Cooperative Perception

### v4

@@ 5.2 Benchmark Results
  • Compared to the results from the Sync setting, the introduction of communication delay in Async settings led to a considerable performance decline. As shown in Tab. 5 and Fig. 7, when subjected to transmission delay with a strict 0.7 IoU threshold, F-Cooper [2], CoAlign [19], and HEAL [18] using LiDAR point cloud experienced significant reductions of 13.30%, 20.49%, and 18.51% respectively. These findings underline the necessity of mitigating the effects of communication delay to ensure effective and robust cooperative perception.
@@
Method M Sync. (AP@IoU = 0.7 / 0.5) $\uparrow$ Async. (AP@IoU = 0.7 / 0.5) $\uparrow$ AM
Overall 0-30m 30-50m 50-100m Overall 0-30m 30-50m 50-100m (MB)
No Fusion L 17.40 / 21.44 18.54 / 22.73 18.60 / 22.22 13.64 / 18.15 17.40 / 21.44 18.54 / 22.73 18.60 / 22.22 13.64 / 18.15 0
Late Fusion L 35.03 / 54.38 52.93 / 67.09 35.46 / 49.56 20.51 / 32.58 22.11 / 47.84 33.13 / 59.97 20.84 / 42.05 14.78 / 28.27 0.003
Early Fusion L 49.04 / 58.89 63.74 / 74.13 43.64 / 52.73 28.71 / 34.38 27.22 / 51.09 35.67 / 64.38 22.05 / 43.26 19.08 / 29.93 1.25
F-Cooper [2] L 45.33 / 60.27 61.87 / 74.91 38.95 / 56.64 21.32 / 35.85 32.03 / 55.56 44.81 / 70.11 23.64 / 50.01 17.31 / 32.48 0.25
CoAlign [19] L 55.56 / 66.42 63.27 / 70.81 54.15 / 66.75 43.01 / 58.36 35.07 / 58.43 39.61 / 63.32 31.83 / 56.64 31.00 / 50.65 0.25
HEAL [18] L 57.78 / 67.75 65.59 / 72.78 56.23 / 69.24 45.32 / 57.79 39.27 / 60.75 43.95 / 65.72 36.04 / 59.99 34.52 / 51.41 0.25
No Fusion C 2.61 / 3.45 1.56 / 5.08 0.29 / 0.69 0.10 / 0.22 2.61 / 3.45 1.56 / 5.08 0.29 / 0.69 0.10 / 0.22 0
Late Fusion C 7.56 / 12.82 16.53 / 27.33 1.93 / 2.38 0.92 / 1.87 4.68 / 11.08 10.19 / 23.88 1.14 / 1.94 0.58 / 1.59 0.003
V2X-ViT [47] C 10.54 / 17.51 19.93 / 32.28 2.64 / 3.42 1.39 / 2.34 7.24 / 15.61 13.58 / 29.19 1.71 / 2.95 0.92 / 2.05 0.25
CoAlign [19] C 12.67 / 19.33 21.05 / 37.01 2.73 / 3.89 1.41 / 2.45 8.61 / 17.33 14.11 / 33.41 1.74 / 3.27 0.97 / 2.17 0.25
HEAL [18] C 13.56 / 19.58 24.84 / 34.74 2.88 / 4.02 1.53 / 2.80 9.53 / 17.42 17.43 / 31.23 1.96 / 3.43 0.98 / 2.47 0.25
No Fusion R 3.06 / 6.98 6.62 / 11.63 0.52 / 1.08 0.26 / 0.87 3.06 / 6.98 6.62 / 11.63 0.52 / 1.08 0.26 / 0.87 0
Late Fusion R 6.87 / 10.95 11.69 / 20.59 1.15 / 1.93 0.54 / 1.21 4.67 / 9.76 7.83 / 18.39 0.74 / 1.44 0.41 / 1.06 0.003
Early Fusion R 7.37 / 13.65 13.89 / 24.07 1.49 / 2.19 0.71 / 1.25 5.24 / 12.09 9.76 / 21.51 1.01 / 1.67 0.46 / 1.09 0.32
V2X-ViT [47] R 8.74 / 14.23 14.92 / 25.23 1.77 / 2.84 0.93 / 1.66 6.01 / 12.68 10.17 / 22.82 1.15 / 2.25 0.61 / 1.46 0.25
CoAlign [19] R 9.76 / 15.59 17.32 / 26.76 1.86 / 3.08 1.23 / 2.05 6.52 / 13.71 11.49 / 23.63 1.16 / 2.41 0.82 / 1.77 0.25
HEAL [18] R 10.08 / 15.82 18.24 / 27.02 1.91 / 3.56 1.12 / 1.83 7.56 / 13.97 12.97 / 24.11 1.23 / 2.85 0.75 / 1.62 0.25
@@ Appendix B Time Synchronization Details
Figure 8: The time synchronization solution for the connected vehicle-side platform and the intelligent roadside unit. Both the vehicle platform and roadside unit are equipped with a computer, a time synchronization box, a GPS/IMU system, and various sensors, including LiDAR, 4D Radar, and cameras. The GPS/IMU systems in both the vehicle platform and roadside unit receive the same GNSS signals from GPS satellites.
@@ Appendix C Robustness Evaluation of Time Synchronization
Table 7: Evaluation of the time synchronization robustness on V2X-Radar-C dataset under varying transmission delays. "M" denotes the modality ("L": LiDAR, "C": Camera, "R": 4D Radar). "AM" indicates the average transmission cost in MB.
@@
No Fusion L 21.44 21.44 21.44 21.44 21.44 0.00
Late Fusion L 54.38 47.84 42.91 40.62 38.66 0.003
F-Cooper [2] L 60.27 55.56 48.10 45.68 44.30 1.25
CoAlign [19] L 66.42 58.43 54.60 51.54 49.82 0.25
HEAL [18] L 67.75 60.75 56.57 53.12 52.17 0.25
No Fusion C 3.45 3.45 3.45 3.45 3.45 0.00
Late Fusion C 12.82 11.08 10.00 9.58 9.12 0.003
F-Cooper [2] C 17.51 15.61 14.01 13.27 12.87 1.25
CoAlign [19] C 19.33 17.33 16.04 15.00 14.50 0.25
HEAL [18] C 19.58 17.42 16.45 15.35 15.08 0.25
No Fusion R 6.98 6.98 6.98 6.98 6.98 0.00
Late Fusion R 10.95 9.76 8.98 8.18 7.79 0.003
F-Cooper [2] R 14.23 12.68 11.81 10.79 10.12 1.25
CoAlign [19] R 15.59 13.71 13.10 12.10 11.08 0.25
HEAL [18] R 15.82 13.97 13.45 12.40 11.25 0.25

### v5

@@ 5.2 Benchmark Results
  • Compared to the results from the Sync setting, the introduction of communication delay in Async settings led to a considerable performance decline. As shown in Tab. 5 and Fig. 7, when subjected to transmission delay with a strict 0.7 IoU threshold, F-Cooper [2], V2X-ViT [47], CoAlign [19], and HEAL [18] using LiDAR point cloud experienced significant reductions. These findings underline the necessity of mitigating the effects of communication delay to ensure effective and robust cooperative perception.
@@
Method M Sync. (AP@IoU = 0.7 / 0.5) $\uparrow$ Async. (AP@IoU = 0.7 / 0.5) $\uparrow$
Overall 0–30m 30–50m 50–100m Overall 0–30m 30–50m 50–100m
No Fusion C 1.00 / 6.76 2.04 / 9.41 0.18 / 5.16 0.01 / 2.27 1.00 / 6.76 2.04 / 9.41 0.18 / 5.16 0.01 / 2.27
Late Fusion C 13.59 / 32.88 16.16 / 40.58 13.29 / 30.37 11.71 / 20.00 9.92 / 30.00 10.88 / 38.64 9.75 / 23.98 8.24 / 15.89
F-Cooper C 15.56 / 44.43 23.22 / 61.97 10.98 / 31.24 4.15 / 15.38 14.25 / 40.90 23.27 / 57.38 6.86 / 29.07 3.62 / 13.74
V2X-ViT C 15.22 / 45.19 20.13 / 57.74 11.20 / 35.31 9.30 / 24.84 13.50 / 43.85 17.24 / 56.28 11.84 / 33.02 8.37 / 20.08
CoAlign C 24.26 / 46.89 36.49 / 63.70 12.75 / 32.77 11.36 / 23.17 21.36 / 45.29 28.86 / 57.47 12.36 / 33.71 11.71 / 22.17
HEAL C 25.05 / 46.94 35.18 / 60.40 13.48 / 33.63 15.80 / 26.88 22.26 / 46.77 32.35 / 60.33 12.55 / 34.12 12.34 / 23.20
No Fusion L 7.13 / 29.09 10.35 / 35.30 4.88 / 25.05 2.52 / 17.72 7.13 / 29.09 10.35 / 35.30 4.88 / 25.05 2.52 / 17.72
Late Fusion L 39.37 / 65.75 50.92 / 80.59 31.59 / 58.64 18.54 / 31.62 34.51 / 62.10 46.85 / 70.32 23.31 / 56.86 15.12 / 23.41
F-Cooper L 50.04 / 73.44 70.29 / 89.52 38.10 / 69.72 17.44 / 34.50 38.99 / 69.38 56.18 / 85.29 27.81 / 63.30 19.65 / 29.90
V2X-ViT L 52.15 / 79.35 68.22 / 88.37 41.79 / 80.74 25.06 / 47.53 39.06 / 73.51 53.71 / 84.06 28.88 / 70.69 15.49 / 42.37
CoAlign L 60.18 / 80.42 75.42 / 91.08 48.10 / 76.48 29.62 / 45.51 53.86 / 77.13 73.65 / 90.58 39.36 / 73.26 17.34 / 35.54
HEAL L 67.57 / 83.00 82.76 / 92.19 57.70 / 80.51 34.79 / 51.93 57.76 / 79.66 74.62 / 90.19 42.33 / 74.18 21.10 / 47.15
No Fusion R 2.69 / 9.02 4.59 / 13.39 0.93 / 5.97 0.22 / 1.28 2.69 / 9.02 4.59 / 13.39 0.93 / 5.97 0.22 / 1.28
Late Fusion R 3.77 / 17.24 6.27 / 25.97 1.44 / 11.19 0.13 / 0.69 3.41 / 15.68 5.26 / 22.85 1.08 / 5.76 0.22 / 2.17
F-Cooper R 6.84 / 23.16 11.80 / 34.98 2.74 / 16.87 0.38 / 2.05 6.37 / 20.70 12.37 / 30.86 2.42 / 12.84 0.16 / 1.74
CoAlign R 11.46 / 26.34 18.01 / 38.46 5.75 / 16.55 0.28 / 2.83 11.12 / 25.49 16.35 / 38.69 3.25 / 14.14 0.22 / 2.34
HEAL R 12.50 / 29.04 23.02 / 44.83 5.16 / 17.85 0.45 / 2.98 11.42 / 25.71 19.41 / 39.30 4.12 / 16.84 0.34 / 2.54
@@ Appendix B Time Synchronization Details
Figure 8: The time synchronization solution for the connected vehicle-side platform and the intelligent roadside unit. Both the vehicle platform and roadside unit are equipped with a computer, a time synchronization box, a GPS/IMU system, and various sensors, including LiDAR, 4D Radar, and cameras. Their GPS/IMU systems receive the same GNSS signals from GPS satellites.
@@ Appendix C Robustness Evaluation of Time Synchronization
Table 7: Evaluation of time synchronization robustness on the V2X-Radar-C dataset under varying transmission delays. Results are reported in AP@IoU=0.5. “M” denotes the sensing modality (L: LiDAR, C: Camera, R: 4D Radar). “AM” indicates the average transmission cost in MB.
@@
No Fusion L 29.09 29.09 29.09 29.09 29.09 0.00
Late Fusion L 65.75 62.10 59.15 53.33 47.76 0.003
F-Cooper [2] L 73.44 69.38 64.15 61.26 55.01 1.25
V2X-ViT L 79.35 73.51 67.23 64.92 61.87 1.25
CoAlign [19] L 80.42 77.13 72.03 68.29 65.02 0.25
HEAL [18] L 83.00 79.66 77.08 73.83 69.52 0.25
No Fusion C 6.76 6.76 6.76 6.76 6.76 0.00
Late Fusion C 32.88 30.00 28.45 24.23 20.64 0.003
F-Cooper [2] C 44.43 40.90 36.72 32.19 26.55 1.25
V2X-ViT C 45.19 43.85 39.14 34.55 31.03 1.25
CoAlign [19] C 46.89 45.29 41.89 37.78 33.35 0.25
HEAL [18] C 46.94 46.77 44.46 39.49 35.13 0.25
No Fusion R 9.02 9.02 9.02 9.02 9.02 0.00
Late Fusion R 17.24 15.68 13.68 12.85 11.56 0.003
F-Cooper [2] R 23.16 20.70 19.52 18.69 16.63 1.25
CoAlign [19] R 26.34 24.49 22.48 21.07 19.12 0.25
HEAL [18] R 29.04 25.71 23.67 22.18 20.53 0.25